OpenStudy (anonymous):

find k if f(x)=e^(x^2-k/x) has a relative min/max at x=2 HELP NEED TO CHECK MY ANSWER FAST THANK YOU

OpenStudy (anonymous):

and your answer is?

OpenStudy (anonymous):

first derivative set to zero subsitute x=2 solve for k voila

OpenStudy (anonymous):

I think that I did the derivative incorrectly now

OpenStudy (anonymous):

chain rule

OpenStudy (anonymous):

I know that e^x = e^x

OpenStudy (anonymous):

right, and?

OpenStudy (anonymous):

I am good at implicit differentiation but the e confuses me

OpenStudy (anonymous):

is it e as is and then the derivative of the exponent?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so I did e^(x^2-k/x) then it is 2x and then I am stuck

OpenStudy (anonymous):

\[\frac{ 1 }{x } = x^{-1}\]

OpenStudy (anonymous):

Sorry. I am lost

OpenStudy (anonymous):

derivative of x^n is \[nx^{n-1}\]

OpenStudy (anonymous):

k is a constant

OpenStudy (anonymous):

ok so that explains the 2x right. The k/x is confusing

OpenStudy (anonymous):

n= -1

OpenStudy (anonymous):

I don't believe it is quotient rule

OpenStudy (anonymous):

k/x = k x^-1

OpenStudy (anonymous):

\[y \prime= e ^{x^2-K/x} \times 2x(kx^-1\] ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

just focus on doing the derivative of this x^2-k/x

OpenStudy (anonymous):

(x^2-k/x)' = 2x -(k/x) '

OpenStudy (anonymous):

k/x = kx^-1

OpenStudy (anonymous):

(kx^-1)' tell me a good reason why you cant apply \[nx^{n-1}\] to that?

OpenStudy (anonymous):

does that turn into kx^-2?

OpenStudy (anonymous):

dont forget the n in front

OpenStudy (anonymous):

-kx^-2

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

ok so then the derivative of x^2-k/x is ????

OpenStudy (anonymous):

2x+kx^-2?

OpenStudy (anonymous):

and kx^-2 is?

OpenStudy (anonymous):

well its k/x^2

OpenStudy (anonymous):

but yes you have it right

OpenStudy (anonymous):

haha I'm understanding this but can you help me put this together

OpenStudy (anonymous):

what do you still not understand?

OpenStudy (anonymous):

so the derivative is 2x+k/x^2?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

derivative of x^2 -k/x is 2x+k/x^2

OpenStudy (anonymous):

what I have now is e^(x^2-k/x) x 2x+k/x^2

OpenStudy (anonymous):

e^(x^2-k/x) (2x+k/x^2) dont use x as a multiplication sign, it can be easily mistaken as a variable use * instead or parenthesis, but yes

OpenStudy (anonymous):

now for the second part of the question. if the relative min/max is at x=2 what Should I do

OpenStudy (anonymous):

relative max/min occurs when f'(x) = 0

OpenStudy (anonymous):

Now like @completeidiot said, we know the derivative is 0 at x = 2 so substitute for x = 2 and set the whole thing to equal 0 and solve for k. @soccerboy43

OpenStudy (anonymous):

hi genius, am i taking too long :P

OpenStudy (anonymous):

na its ok. take ur time. u still got 257 days, 3 hours, and 27 seconds.

OpenStudy (anonymous):

plug 2 in for all x values?

OpenStudy (anonymous):

till?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

till this thing ends.

OpenStudy (anonymous):

made up number?

OpenStudy (anonymous):

ya lawl. i think u missed the joke there...

OpenStudy (anonymous):

i think i did

OpenStudy (anonymous):

I have e^(4-k/4) times 4 + (2/x)^2

OpenStudy (anonymous):

you forgot an x

OpenStudy (anonymous):